How would I solve
2x-3y=6???

Answer by checkley71(8403)

2X-3Y=6 IF THIS IS AN EQUATION FOR A LINE THEN
-3Y=-2X+6
Y=-2X/-3+6/-3
Y=2X/3-2 THIS IS A LINE WITH A SLOPE OF 2/3 AND A Y INTERCEPT OF -2 (0,-2)
SETTING Y=0 WE CAN FIND THE X INTERCEPT
0=2X/3-2
2X/3=2
2X=6
X=6/2
X=3 OR (3,0)